# Balancing Act: Prioritization Strategies for LLM-Designed Restless Bandit Rewards

### v4

@@ 5.2.3 Error Bounds for Adjudicator’s Selection
Relative $\textit{Relative Regret}=\frac{f_{p}(\mathbf{s}(R^{*}))-f_{p}(\mathbf{s}(\tilde{R}^{*}))}{f_{p}(\mathbf{s}(R^{*}))}$ (3)
@@ 6 Experiments
We describe our testset (Section 6.1), the compared methods (Section 6.2), and our experimental results both for dealing with composite prioritization prompts (Section 6.3) and additionally minimizing unintended side effects (Section 6.4). Following the work of Behari et al. [7], which constitutes our most important baseline, we use Gemini Pro [2] as the LLM in our experiments.
@@ 6.1 Dataset Description
Instances of our problem consist of two parts: A preference prompt and an RMAB instance. We initially focus on prioritization prompts. Specifically, for each feature $X$, we consider two different prioritization clauses “Prioritize agents with low/high value of feature $X$”. This gives rise to $6$ singular prompts consisting of one prioritization clause, two for each feature. For composite prompts, we take all combinations of two features and the two prioritization clauses for each feature (e.g. “Prioritize agents with high value of feature $A$ and also prioritize agents with low value of feature $B$”). This results in $3\cdot 4=12$ composite prompts. For each domain, we run each prompt on the $15$ RMAB instances from the three datasets.
@@ 6.2 Models & Baselines
We analyze six different variants of SCLM differing in the used social welfare function (Utilitarian, Egalitarian, Nash) and scorer model (Simulator or LLM), e.g., we denote as SCLM-SIM-Egalitarian SCLM with the Simulator Scorer Model and the Egalitarian social welfare function. In our generator, we generate $4$ candidate reward functions in each step and run $5$ iterations to generate a total of $20$ candidate reward functions. In addition, we consider several LLM-focused baselines (see Appendix F for detailed descriptions):
@@
This baseline implements the Decision-Language Model by Behari et al. [7] (see Section C.2).
@@
This is a modified version of DLM where within the reflection prompt, we include examples for singular queries of how the LLM should reason over the different reward function choices (see Appendix F).
@@ 6.3 Results: Composite Prioritization Prompts
We analyze the performance on the $12$ composite prompts described above which request the prioritization of two subpopulations (see Section E.1 for additional results).
@@
[…]
@@
Figure 4: Results comparing the quality of reward design methods for composite prioritization prompts. Results are averaged across $180=12\cdot 15$ values: $12$ composite prompts on $15$ RMAB instances (from $3$ datasets). Error bars represent std-error.
@@ Results
In Figure 4, we show the averaged results from the synthetic and real-world domain. We depict the average summed and minimum alignment with the two clauses of the composite prompt, i.e., the minimum/summed change in the utility generated by the prioritized group of agents.
[…]
@@
DLM-PrioritizationOnly 6.809$\pm$0.86 0.302$\pm$0.02
DLM-ExtendedPrompt-Fair -0.254$\pm$0.73 0.276$\pm$0.02
SCLM-PrioritizationOnly 13.131$\pm$0.86 0.316$\pm$0.02
SCLM-ExtendedPrompt-Fair 15.364$\pm$0.94 0.099$\pm$0.01
@@ 6.4 Addressing Fairness and Biases
As discussed in Section 5.2, we can also use our pipeline to prevent unintended side-effects of aligning reward functions with prioritization clauses, i.e., (i) shifts in the utility feature distribution of features not included in the prompt, (ii) drops in the total generated utility, and (iii) arbitrary weighted combinations of the above two goals and additional preference clauses. We focus on (i) here and relegate the results for (ii) and (iii), which paint a very similar picture, to Appendix D.3.
We analyze all $6$ singular and $12$ composite prioritization prompts (see Section 6.1), where we add additional clauses to prevent shifts in the utility distribution of all features not referenced in the prompt. We use the simulator scorer with a Utilitarian social welfare function and call the resulting model SCLM-ExtendedPrompt-Fair. As baselines, we consider DLM only prompted with the prioritization clause(s) (called DLM-PrioritizationOnly) and DLM prompted with the prioritization clause(s) and clause(s) together for a request for minimizing of utility shifts for the other features (called DLM-ExtendedPrompt-Fair). We also consider SCLM-SIM-Utilaterian only prompted with the prioritization clause(s) (called SCLM-PrioritizationOnly). See Appendix E for more details on the prompts.
To compute the alignment with prioritization clauses, similar to Section 6.3, we compute average change in utility generated by prioritized subpopulations. To quantify unintended utility shifts, we compute the average Earth mover’s distance between utility feature distribution under the candidate and default reward function for each feature not included in one of the prioritization clauses.
Table 1 shows the results. Comparing DLM-PrioritizationOnly and DLM-ExtendedPrompt-Fair, we find that adding additional objective to the prompt does not result in a better performance for real-world domains. In contrast, SCLM-ExtendedPrompt-Fair which incorporates unintended shifts in the selection chooses reward functions resulting in significantly higher utility increases for prioritized subpopulations and significantly fewer unintended utility shifts. The fact that SCLM performs advantageously for both (conflicting) objectives highlights the quality of the pipeline and its capabilities to effectively address multiple objectives (of different types). We see similar results in synthetic domain (see Table 5 in the appendix).
@@ Appendix A Additional Related Work
Appendix B Additional Material for Section 4
@@
Moving from the default reward function to a reward function aligned with a given (prioritization) prompt causes shifts within the distribution of resources and utility. Due to correlations between features, this might change might lead to unintended utility shifts for features not specified in the prompt. In Figure 5, we show an instance of our experiments. We present the utility feature distribution for the two features income and education for two reward functions: The reward function selected by DLM for the prompt “ Prioritize agents with low income ” (orange) and the default reward (blue). While the utility generated by low-income agents increases, the utility generated by highly educated agents decreases, a side-effect a user is likely unaware of and that might conflict with their allocation goals.
@@ Appendix D Error Bound in Adjudicator
$f_{p}(\mathbf{s}(\tilde{R}^{*}))\geq\alpha\cdot f_{p}(\tilde{\mathbf{s}}(\tilde{R}^{*})).$ (14)
@@
Relative $\textit{Relative Regret}=\frac{f_{p}(\mathbf{s}(R^{*}))-f_{p}(\mathbf{s}(\tilde{R}^{*}))}{f_{p}(\mathbf{s}(R^{*}))}\leq 1-\alpha^{2}$ (18)
@@ E.1 Composite Prioritization Prompts
DLM-PrioritizationOnly 6.809$\pm$0.86 0.302$\pm$0.02 0.06$\pm$1.13 1.15$\pm$0.04
DLM-ExtendedPrompt-Fair -0.254$\pm$0.73 0.276$\pm$0.02 -1.882$\pm$1.31 1.173$\pm$0.06
SCLM-PrioritizationOnly 13.131$\pm$0.86 0.316$\pm$0.02 9.349$\pm$1.45 1.218$\pm$0.05
SCLM-ExtendedPrompt-Fair 15.364$\pm$0.94 0.099$\pm$0.01 10.006$\pm$1.13 0.239$\pm$0.02
@@
DLM-PrioritizationOnly 6.809$\pm$0.86 0.317$\pm$0.01 -6.233$\pm$0.17 0.06$\pm$1.13 0.485$\pm$0.01 -5.403$\pm$0.28
DLM-ExtendedPrompt-Util 3.416$\pm$1.04 0.347$\pm$0.02 -5.689$\pm$0.25 -2.508$\pm$1.2 0.548$\pm$0.02 -3.825$\pm$0.29
SCLM-PrioritizationOnly 13.131$\pm$0.86 0.397$\pm$0.01 -5.372$\pm$0.19 9.349$\pm$1.45 0.614$\pm$0.01 -2.8$\pm$0.24
SCLM-ExtendedPrompt-Util 8.869$\pm$0.73 0.519$\pm$0.02 -3.91$\pm$0.21 13.211$\pm$1.31 0.736$\pm$0.01 -0.568$\pm$0.12
SCLM-Full 11.496$\pm$0.87 0.529$\pm$0.01 -3.85$\pm$0.18 9.166$\pm$1.05 0.705$\pm$0.01 -1.044$\pm$0.11
@@ E.2 Effect of Social Choice Function
Figure 6: Comparison of different methods using the prompt “Prioritize agents with a low value for feature $B$ and agents with a low value for feature $C$”.
While both the Utilitarian and Egalitarian social welfare function leads to good overall results, the choice between them significantly influences which reward function is selected at the instance level. Figure 6 shows one example instance for a prompt with two prioritization clauses. As in Figure 1, each point corresponds to a candidate reward function with the axes measuring alignment with the two respective clauses. (as described in the beginning of the section 6.3). The SCLM-Utilitarian model chooses a reward function from the Pareto frontier that shows a much stronger effect for the second clause (i.e., the utility increase for agents with a low value of feature $C$ is more pronounced). In contrast, the reward function selected by SCLM-Egalitarian is much more balanced. DLM-based baselines fail to pick a reward function from the Pareto frontier.
@@ E.3 Preventing Ineffective Resource Use
We additionally consider situations where the user specifies prioritization prompts for multiple attributes but is still interested in maximizing the total generated utility. As discussed at the end of Section 5.2, our framework can handle such settings by adding utility maximization as an additional evaluation dimension. For all 6 singular and 12 composite prioritization prompts in our tests, we add additional clauses to maximize total utility and call the resulting model SCLM-ExtendedPrompt-Util. As a baseline, we consider three models: DLM/SCLM prompted only by the prioritization clause (DLM/SCLM-Prioritization Only) and DLM prompted by both the prioritization clause(s) and a clause requesting the maximization of overall utility (DLM-ExtendedPrompt-Util).
@@
Finally, we consider a variant of SCLM which both prevents unintended shifts and utility drops, i.e., in addition to the specified prioritization clauses, we add clause(s) on the minimization of shifts and a clause on the minimization of utility drops (which results in having a total of 4 evaluation dimensions; one per feature plus the utility drop one). We use the simulator scorer together with the utilitarian welfare function and call the resulting model SCLM-Full. We see in the last row of Table 6 that this model chooses balanced reward functions with high % change in utility in desired attributes, high overall utility, and low unintended shift, showcasing that SCLM can flexibly allow the end-user to combine multiple evaluation dimensions.

### v5

@@ 5.2.3 Error Bounds for Adjudicator’s Selection
$\textit{Relative Regret}=\frac{f_{p}(\mathbf{s}(R^{*}))-f_{p}(\mathbf{s}(\tilde{R}^{*}))}{f_{p}(\mathbf{s}(R^{*}))}$ (3)
@@ 6 Experiments
We describe our testset (Section˜6.1), the compared methods (Section˜6.2), and our experimental results both for dealing with composite prioritization prompts (Section˜6.3) and additionally minimizing unintended side effects (Section˜6.4). Following the work of Behari et al. [7], which constitutes our most important baseline, we use Gemini Pro [2] as the LLM in our experiments.
@@ 6.1 Dataset Description
Instances of our problem consist of two parts: A preference prompt and an RMAB instance. We initially focus on prioritization prompts. Specifically, for each feature $X$, we consider two different prioritization clauses “Prioritize agents with low/high value of feature XX”. This gives rise to $6$ singular prompts consisting of one prioritization clause, two for each feature. For composite prompts, we take all combinations of two features and the two prioritization clauses for each feature (e.g. “Prioritize agents with high value of feature $A$ and also prioritize agents with low value of feature BB”). This results in $3\cdot 4=12$ composite prompts. For each domain, we run each prompt on the $15$ RMAB instances from the three datasets.
@@ 6.2 Models & Baselines
We analyze six different variants of SCLM differing in the used social welfare function (Utilitarian, Egalitarian, Nash) and scorer model (Simulator or LLM), e.g., we denote as SCLM-SIM-Egalitarian SCLM with the Simulator Scorer Model and the Egalitarian social welfare function. In our generator, we generate $4$ candidate reward functions in each step and run $5$ iterations to generate a total of $20$ candidate reward functions. In addition, we consider several LLM-focused baselines (see Appendix˜F for detailed descriptions):
@@
This baseline implements the Decision-Language Model by Behari et al. [7] (see Section˜C.2).
@@
This is a modified version of DLM where within the reflection prompt, we include examples for singular queries of how the LLM should reason over the different reward function choices (see Appendix˜F).
@@ 6.3 Results: Composite Prioritization Prompts
We analyze the performance on the $12$ composite prompts described above which request the prioritization of two subpopulations (see Section˜E.1 for additional results).
@@
Instead, we assume that the terms “low” and “high” in the input prompts refer to the most extreme feature values. Let $p_{i}$ be some prompt prioritizing agents with a high/low value of some feature $X$. As the evaluation score $e_{i}(R)$, we compute the summed utility generated by the agents with highest/lowest value of $X$ under the policy $\Pi(R)$ normalized by the utility generated by these agents under the default policy $\Pi(R^{*})$. [^2] In Section E.1, we also check how the results change if we instead interpret “low”/“high” to refer to the lowest/highest two or three values. We observe very similar trends. Reflecting the multi-objective nature of our problem, we consider two metrics for measuring the alignment of a reward $R$ with a full composite prompt: sum and minimum of $\%$ change of the utility generated by the two prioritized groups under policy $\Pi(R)$ compared to the default policy, i.e., the sum (resp. minimum) of the evaluation scores for $R$.
@@
Figure 4: Results comparing the quality of reward design methods for composite prioritization prompts. Results are averaged across $180=12\cdot 15$ values: 1212 composite prompts on 1515 RMAB instances (from 33 datasets). Error bars represent std-error.
@@ Results
In Figure˜4, we show the averaged results from the synthetic and real-world domain. We depict the average summed and minimum alignment with the two clauses of the composite prompt, i.e., the minimum/summed change in the utility generated by the prioritized group of agents.
We start by focusing on SCLM with Simulator Scorer SCLM-SIM (green-shaded bars), our strongest method. On both domains, SCLM-SIM significantly outperforms all baselines for both minimum and summed $\%$ change independent of whether the Utilitarian, Egalitarian or Nash social welfare function is chosen. SCLM-SIM-Utilitarian outperforming the baselines for the minimum change and SCLM-SIM-Egalitarian outperforming them for the summed change highlights the advantages of the SCLM, as these objectives are not explicitly optimized by the respective models, e.g., SCLM-SIM-Utilitarian aims at maximizing the summed change and not the minimum one, but still performs well regarding the minimum change. This indicates that SCLM independent of the chosen welfare function does a better job at picking effective and aligned reward functions (on the Pareto front). Comparing SCLM-SIM-Utilitarian and SCLM-SIM-Egalitarian, the two methods exhibit a big difference under the summed change criterion, while the difference regarding the minimum change is much smaller. Examining individual instances we find in Section˜E.1 that both functions lead to very different selections on the instance level; unsurprisingly, the Egalitarian method creates rewards that benefit both groups in a more balanced fashion. For the Nash welfare function, we found that the performance was similar, yet slightly inferior to the Utilitarian welfare function in all relevant evaluation dimensions.
@@
[…]
@@ 6.4 Addressing Fairness and Biases
As discussed in Section˜5.2, we can also use our pipeline to prevent unintended side-effects of aligning reward functions with prioritization clauses, i.e., (i) shifts in the utility feature distribution of features not included in the prompt, (ii) drops in the total generated utility, and (iii) arbitrary weighted combinations of the above two goals and additional preference clauses. We focus on (i) here and relegate the results for (ii) and (iii), which paint a very similar picture, to Appendix D.3.
We analyze all $6$ singular and $12$ composite prioritization prompts (see Section˜6.1), where we add additional clauses to prevent shifts in the utility distribution of all features not referenced in the prompt. We use the simulator scorer with a Utilitarian social welfare function and call the resulting model SCLM-ExtendedPrompt-Fair. As baselines, we consider DLM only prompted with the prioritization clause(s) (called DLM-PrioritizationOnly) and DLM prompted with the prioritization clause(s) and clause(s) together for a request for minimizing of utility shifts for the other features (called DLM-ExtendedPrompt-Fair). We also consider SCLM-SIM-Utilaterian only prompted with the prioritization clause(s) (called SCLM-PrioritizationOnly). See Appendix E for more details on the prompts.
To compute the alignment with prioritization clauses, similar to Section˜6.3, we compute average change in utility generated by prioritized subpopulations. To quantify unintended utility shifts, we compute the average Earth mover’s distance between utility feature distribution under the candidate and default reward function for each feature not included in one of the prioritization clauses.
Table˜1 shows the results. Comparing DLM-PrioritizationOnly and DLM-ExtendedPrompt-Fair, we find that adding additional objective to the prompt does not result in a better performance for real-world domains. In contrast, SCLM-ExtendedPrompt-Fair which incorporates unintended shifts in the selection chooses reward functions resulting in significantly higher utility increases for prioritized subpopulations and significantly fewer unintended utility shifts. The fact that SCLM performs advantageously for both (conflicting) objectives highlights the quality of the pipeline and its capabilities to effectively address multiple objectives (of different types). We see similar results in synthetic domain (see Table 5 in the appendix).
@@ Appendix A Additional Related Work
Appendix B Additional Material for Section˜4
@@
Moving from the default reward function to a reward function aligned with a given (prioritization) prompt causes shifts within the distribution of resources and utility. Due to correlations between features, this might change might lead to unintended utility shifts for features not specified in the prompt. In Figure˜5, we show an instance of our experiments. We present the utility feature distribution for the two features income and education for two reward functions: The reward function selected by DLM for the prompt “ Prioritize agents with low income ” (orange) and the default reward (blue). While the utility generated by low-income agents increases, the utility generated by highly educated agents decreases, a side-effect a user is likely unaware of and that might conflict with their allocation goals.
@@ Appendix D Error Bound in Adjudicator
$).f_{p}(\mathbf{s}(\tilde{R}^{*}))\geq\alpha\cdot f_{p}(\tilde{\mathbf{s}}(\tilde{R}^{*}$)). (14)
@@
$\textit{Relative Regret}=\frac{f_{p}(\mathbf{s}(R^{*}))-f_{p}(\mathbf{s}(\tilde{R}^{*}))}{f_{p}(\mathbf{s}(R^{*}))}\leq 1-\alpha^{2}$ (18)
@@ E.1 Composite Prioritization Prompts
[…]
@@
[…]
@@ E.2 Effect of Social Choice Function
Figure 6: Comparison of different methods using the prompt “Prioritize agents with a low value for feature $B$ and agents with a low value for feature CC”.
While both the Utilitarian and Egalitarian social welfare function leads to good overall results, the choice between them significantly influences which reward function is selected at the instance level. Figure˜6 shows one example instance for a prompt with two prioritization clauses. As in Figure 1, each point corresponds to a candidate reward function with the axes measuring alignment with the two respective clauses. (as described in the beginning of the section 6.3). The SCLM-Utilitarian model chooses a reward function from the Pareto frontier that shows a much stronger effect for the second clause (i.e., the utility increase for agents with a low value of feature $C$ is more pronounced). In contrast, the reward function selected by SCLM-Egalitarian is much more balanced. DLM-based baselines fail to pick a reward function from the Pareto frontier.
@@ E.3 Preventing Ineffective Resource Use
We additionally consider situations where the user specifies prioritization prompts for multiple attributes but is still interested in maximizing the total generated utility. As discussed at the end of Section˜5.2, our framework can handle such settings by adding utility maximization as an additional evaluation dimension. For all 6 singular and 12 composite prioritization prompts in our tests, we add additional clauses to maximize total utility and call the resulting model SCLM-ExtendedPrompt-Util. As a baseline, we consider three models: DLM/SCLM prompted only by the prioritization clause (DLM/SCLM-Prioritization Only) and DLM prompted by both the prioritization clause(s) and a clause requesting the maximization of overall utility (DLM-ExtendedPrompt-Util).
@@
Finally, we consider a variant of SCLM which both prevents unintended shifts and utility drops, i.e., in addition to the specified prioritization clauses, we add clause(s) on the minimization of shifts and a clause on the minimization of utility drops (which results in having a total of 4 evaluation dimensions; one per feature plus the utility drop one). We use the simulator scorer together with the utilitarian welfare function and call the resulting model SCLM-Full. We see in the last row of Table˜6 that this model chooses balanced reward functions with high % change in utility in desired attributes, high overall utility, and low unintended shift, showcasing that SCLM can flexibly allow the end-user to combine multiple evaluation dimensions.